OpenStudy (anonymous):

Another quantity recorded in the aforementioned research paper is the toughness of the spider silk, found to be 27 J/cm^3 . Toughness is the amount of energy that an object can absorb, per unit volume of the object, before breaking. Spiderman and Mary Jane freefall a height 66.9m before having their fall halted by a strand of web 1.00cm in diameter fired upwards a height 66.9m (so the strand of web is long). If all of their kinetic energy is absorbed by the spider silk, what is their maximum combined mass?

OpenStudy (anonymous):

I think I understand the question. Our prof. has taught us about Tensile strength so I'm sure we're supposed to apply it here. Tensile strength = F/A What I did was change 27J/cm^3 to N/m^2 to get 2.7x10^7N/m^2. Since they give you the length and width of the web, I converted 1.00cm to 0.01m and multiplied that by 66.9m, to get 0.669m^2. They ask for the mass, and we've learned that in this case F = mg. Therefore I get an equation that looks something like this: 2.7x10^7N/m^2 = m(9.81)/0.669 Now I should be able to solve this easily but this gives me a huge mass which is physically impossible. I was wondering if anyone would be able to show me my error. Thanks alot.

OpenStudy (asnaseer):

the spider wen strand is 1.00cm in "diameter". so work out the volume of the strand you need to use area of circle formulae and then multiply by length of strand. hope that helps.

OpenStudy (asnaseer):

wen == web

OpenStudy (anonymous):

But how do you find the area from the volume? I dont really understand what you were trying to say there.

OpenStudy (asnaseer):

remember the question stated the toughness of the spider silk to be "27 J/cm^3" - i.e. it is based on volume not area.